AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Aug 11 09:25:13 2022
Author
*
:
Subject
*
:
<p><br /> Moved pneumatic motor in and out several revolutions</p> <p>Biased detector up to 100 V, I ~ 12.4 uA ambient temp c. 21 deg C</p> <p>Moving detector in increased leakge current to ~14.5 uA, moving detector out dropped leakage current back to ~12.4 uA</p> <p>aida05 - pulser width 21 chn (slightly lower than yesterday)</p> <p>aida01 - pulser width 124 chn</p> <p>aida02 - double peak (is not currently grounded)</p> <p>aida03 - 183 chn</p> <p>aida04 - 212 chn</p> <p> </p> <p>Attachment 1 1.8.L spectra p+n FEE64s</p> <p>Attachment 2 per FEE64 rate spectra</p> <p>Attachment 3 aida05 1.8.L, 1.8.W and rate spectra</p> <p>Attachment 4 FEE64 temperatures - OK</p> <p>Attachment 5 per FEE64 ADC data event stats - rates for aida01-aida04 significabtly lower than yesterday presumably due to detector bias -100V</p> <p>Attachments 6 & 7 per FEE64 1.8.W spectra</p> <p>Attachment 8 1.8.L spectra for n+n FEE64s</p> <p>regrounded aida02 back to cryring ground - required removing star configuration of all fee grounds and making a new</p> <p>aida05 - pulser width 30 chn </p> <p>aida01 - pulser width 300 chn</p> <p>aida02 - 220 -> double peak gone </p> <p>aida03- 190</p> <p>aida04 - 200</p> <p>Perhaps star configuration isn't as tight as before - re tighten star section</p> <p>Re tightened section -> no impact on pulser widths</p> <p>Moved detector in using pneumatic -> no impact on pulser widths</p> <p>Added a grounding cable from the 19 inch rack to the same CRYRING ground that the fees are connected too. Stats around 2-3k, slow comparator = 0x64, bias=100v, I=12.18 uA</p> <p>aida01 ~ 260</p> <p>aida02 ~ 177</p> <p>aida03 ~ 178</p> <p>aida04 ~ 187</p> <p>aida05 ~28</p> <p>Removed CRYRING ground. Fees and rack connected together still. </p> <p>no change</p> <p>Bias supply module is floating</p> <p>ion pumps off, gauge off -> no change in waveforms, leakage current dropped to 8.7 uA</p> <p>aida01 ~ 255</p> <p>aida02 ~ 176</p> <p>aida05 ~ 26</p> <p>Removed all vacuum systems and motors power from 19 inch rack power. Connected to the rack are now: NIM bin, network switch, 2 inputs to mains relay, raspberrypi, pi dlink usb connector, VME crack</p> <p>aida01 ~ 220</p> <p>aida02 ~ 154</p> <p>aida05 ~ 27</p> <p>aida03 ~ 181</p> <p>aida04 ~ 193</p> <p>Moved power cable for 19 inch rack to CRYRING 'clean' power source</p> <p>aida03 ~ 211</p> <p>aida04 ~ 212</p> <p>aida01 ~ 317</p> <p>aida02 ~ 215</p> <p>aida05 ~ 29</p> <p>Reconnected CRYRING ground </p> <p>aida01 ~ 301</p> <p>aida02 ~ 203</p> <p>aida05 ~ 28</p> <p>aida03 ~ 181</p> <p>aida04 ~ 191</p> <p>Current status 17:00:<br /> All CARME vacuum equipment is off (turbo, SIP power modules, NEG power modules, gauges) - roughing pump is on but is valved off from the turbo which is valved off from chamber</p> <p>Motors and motors pi are off. Detectors are fully out and the pin has been reinserted. </p> <p>Bias ON 100V, 8.5 uA. Pulser ON, all equipment connected to 19 inch rack power ON. Jumper is on link 1 for nn bias adaptor card only. </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
19.PNG
Original size: 1920x1031
Attachment 2:
18.PNG
Original size: 1920x1031
Attachment 3:
17.PNG
Original size: 1920x1031
Attachment 4:
16.PNG
Original size: 1920x1031
Attachment 5:
15.PNG
Original size: 1920x1031
Attachment 6:
14.PNG
Original size: 1920x1031
Attachment 7:
13.PNG
Original size: 1920x1031
Attachment 8:
12.PNG
Original size: 1920x1031
Attachment 9:
11.PNG
Original size: 1920x1031
Attachment 10:
10.PNG
Original size: 1920x1031
Attachment 11:
9.PNG
Original size: 1920x1031
Attachment 12:
8.PNG
Original size: 1920x1031
Attachment 13:
7.PNG
Original size: 1920x1031
Attachment 14:
6.PNG
Original size: 1920x1031
Attachment 15:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown